### v3

# Translating predictive distributions into informative priors

Abstract
When complex Bayesian models exhibit implausible behaviour, one solution is to assemble available information into an informative prior. Challenges arise as prior information is often only available for the observable quantity, or some model-derived marginal quantity, rather than directly pertaining to the natural parameters in our model. We propose a method for translating available prior information, in the form of an elicited distribution for the observable or model-derived marginal quantity, into an informative joint prior. Our approach proceeds given a parametric class of prior distributions with as yet undetermined hyperparameters, and minimises the difference between the supplied elicited distribution and corresponding prior predictive distribution. We employ a global, multi-stage Bayesian optimisation procedure to locate optimal values for the hyperparameters. Three examples illustrate our approach: a cure-fraction survival model, where censoring implies that the observable quantity is a priori a mixed discrete/continuous quantity; a setting in which prior information pertains to $R^{2}$ – a model-derived quantity; and a nonlinear regression model.
Keywords: prior specification, multi-objective optimisation, prior predictive checks
1 Introduction
Incorporating prior information in Bayesian models is conceptually easy, but in practice constructing an informative prior is not easy. Formulating priors in accordance with predictive information obtained via predictive elicitation (O’Hagan et al., 2006) is attractive due to the widespread availability, reliability (Kadane & Wolfson, 1998) and model-agnostic nature of such information. However it is often unclear how to implement this approach, particularly for complex, nonlinear, or overparameterised models, for which informative priors can be essential to exclude model behaviours that conflict with known properties of the world. In this paper we suppose predictive information is available in the form of a target prior predictive distribution, and consider how to translate this into a prior distribution for model parameters, a step that has heretofore received relatively little attention.
One simple approach to this task is to directly model the elicited quantity. This requires no translation step. For example, Perepolkin et al. (2021) directly updated elicited information in light of observations using a Bayesian quantile-parameterised likelihood. Such direct approaches are currently only feasible for simple models with no latent structure. For models with simple latent structure, eliciting information about an invertible function of the parameters may be possible (e.g. Chaloner et al., 1993), enabling analytic translation into a prior for the parameters. Translation is also clear for conjugate distributions, since the prior predictive distribution determines the prior hyperparameter values (Percy, 2002). Translation, however, is unclear in general for nonconjugate models (Gribok et al., 2004). Techniques exist for specific models with specific latent structures, including for logistic regression (Chen et al., 1999), contingency table analyses (Good, 1967) and hierarchical models (Hem, 2021), but a model-agnostic approach is needed for models outwith these classes, as noted by Mikkola et al. (2023).
Our approach to translation builds on the idea of predictive checks Gabry et al., 2019; Winkler, 1967, the “hypothetical future samples” of and the Bayesian workflow (Gelman et al., 2020), in which the prior is repeatedly adjusted until there is concordance between the prior predictive distribution and the elicited predictive information. However, this manual approach is impractical whenever the relationship between the prior and the distribution of the observables is muddied by the complexity of the intervening model. A more automated method is required. Wang et al. (2018) and Thomas et al. (2020) have proposed approaches in which either regions of observable space or specific realisations are labelled as plausible or implausible by experts, and then a prior accounting for this information is formed via either history matching or a “human in the loop” process driven by a Gaussian process model. Albert et al. (2012) propose a supra-Bayesian approach intended for multiple experts, in which a Bayesian model is formed for quantiles or probabilities elicited from the experts. Another approach, and the closest in motivation and methodology to ours, is Hartmann et al. 2020; da Silva et al., 2019, which is partly inspired by, which employs a Dirichlet likelihood for elicited predictive quantiles to handle both elicitation and translation. Our approach is model-agnostic and is fully based around distributions, meaning uncertainty is directly and intuitively represented. We specify a suitable, generic loss function between the prior predictive distribution and this target distribution, and minimise this loss function via a generic, multi-objective global optimisation process. We implement our methodology in an R package pbbo (https://github.com/hhau/pbbo; Supplement S1).
2 Methodology
We postulate three properties that we would like our method to satisfy:
Faithfulness   A prior is faithful if it accurately encodes the target data distribution provided by the elicitation subject. Faithfulness is a property of both the procedure employed to obtain the prior and the model itself, since not all target prior predictive distributions can be encoded by simple models and prior structures. Faithfulness is related to the definition of validity in Johnson et al. (2010) and O’Hagan et al. (2006)’s use of faithful, but their concerns are specific to the elicitation process and not to the translational step.
Uniqueness   Multiple equally faithful prior distributions may exist in complex models, meaning we must distinguish between such priors based on properties other than just faithfulness if a unique prior is desired. One approach is to choose the prior with the largest marginal standard deviations (see Section 2.3), but other properties could be used similarly. The challenge of uniqueness has been noted by Stefan et al. (2022).
Replicability   A procedure is replicable if, given the same target, it constructs identical priors across independent replications. This is unlikely to hold exactly with stochastic algorithms, meaning it is important to assess.
2.1 The target predictive distribution $\text{T}(Y)$
We assume a target predictive distribution, with cumulative distribution function (CDF) $\text{T}(Y)$, for the observable quantity $Y\in\mathcal{Y}\subseteq{}\mathbb{R}$ has been chosen. We recommend chosing this using predictive elicitation (Kadane & Wolfson, 1998), in which experts are queried about the observable quantity at a small number quantiles, then an appropriate parametric distribution is fitted to the elicited values (O’Hagan et al., 2006, chap. 6). However, our approach applies regardless of how $\text{T}(Y)$ is selected, although we assume it can be described by a (mixture of) standard distributions and that samples can be drawn from it.
We describe our methodology in the more general setting in which we wish to elicit information about the observable quantity $Y$ conditional on some known values of a covariate $X\in\mathcal{X}\subseteq\mathbb{R}^{C}$. For example, when using the linear model $Y=X\beta+\varepsilon$ we may elicit information about $Y$ at a fixed set of values for $X$ corresponding to an experimental design. Specifically we suppose the target CDF $\text{T}(Y\mid X_{r})$ has been elicited at $R$ values of the covariate vector denoted $\{X_{r}\}_{r=1}^{R}$, which we stack in the matrix $\boldsymbol{X}=\left[X_{1}^{\top}\cdots X_{R}^{\top}\right]\in\boldsymbol{\mathcal{X}}\subseteq\mathbb{R}^{R\times C}$.
2.2 Predictive discrepancy (primary objective)
Consider a joint model for observables $Y$ and parameters $\theta\in\Theta\subseteq\mathbb{R}^{Q}$, given hyperparameters $\lambda\in\Lambda\subset\mathbb{R}^{L}$ and covariates $X$. This joint model has CDF $\text{P}(Y,\theta\mid\lambda,X)$ and prior predictive CDF $\text{P}(Y\mid\lambda,X)$ for $Y$. We assume that each target $\text{T}(Y\mid X_{r})$ has identical support to $\text{P}(Y\mid\lambda,X_{r})$. We denote $\text{T}(Y\mid\boldsymbol{X})=\prod_{r=1}^{R}\text{T}(Y\mid X_{r})$, with $\text{P}(Y\mid\lambda,\boldsymbol{X})$ and $\text{P}(\theta\mid\lambda,\boldsymbol{X})$ defined analogously.
We quantify the difference between the prior predictive and target by the covariate-specific predictive discrepancy, which we define to be
$\tilde{D}(\lambda\mid\boldsymbol{X})=\frac{1}{R}\sum_{r=1}^{R}\int d(\text{P}(Y\mid\lambda,X_{r}),\text{T}(Y\mid X_{r}))\text{d}\text{T}(Y\mid X_{r}),$ (1)
for some discrepancy function $d(\cdot,\cdot)$. Minimising (1) admits the optimal hyperparameter $\lambda^{*}=\min_{\lambda\in\Lambda}\tilde{D}(\lambda\mid\boldsymbol{X})$. The covariate-independent equivalent $\tilde{D}(\lambda)$ is obtained by setting $R=1$ and ignoring conditioning on $X_{r}$.
Inspired by the Cramér-von Mises (von Mises, 1947) and Anderson-Darling (Anderson & Darling, 1952) distributional tests we define, for CDFs $\text{M}(Y)$ and $\text{P}(Y)$, two options for the discrepancy function,
$d^{\text{CvM}}(\text{M}(Y),\text{P}(Y))=\left(\text{M}(Y)-\text{P}(Y\right))^{2},\quad d^{\text{AD}}(\text{M}(Y),\text{P}(Y))=\frac{\left(\text{M}(Y)-\text{P}(Y\right))^{2}}{\text{P}(Y)(1-\text{P}(Y))}.$ (2)
Both discrepancies are proper scoring rules (Gneiting & Raftery, 2007) as they are minimised iff $\text{M}(Y)=\text{P}(Y)$ for all $Y\in\mathcal{Y}$. Supposing $\text{P}(Y\mid\lambda,X_{r})$ is flexible enough to exactly match $\text{T}(Y\mid X_{r})$ for some $\lambda^{*}$, then both discrepancies will yield the same $\lambda^{*}$. Differences arise when $\text{P}(Y\mid\lambda,X_{r})$ is insufficiently flexible, in which case the Anderson-Darling (AD) discrepancy $d^{\text{AD}}$ places more emphasis on matching the tails of two CDFs.
2.3 Regularising estimates of $\lambda^{*}$ (secondary objective)
There often are many optimal values $\lambda^{*}$ that yield values of $\tilde{D}(\lambda^{*}\mid\boldsymbol{X})$ that are practically indistinguishable (noted by da Silva et al., 2019) but with immensly differing prior distributions $\text{P}(\theta\mid\lambda^{*},\boldsymbol{X})$. That is, there are many equally faithful priors. This is not surprising because we are providing information only on $Y$, which is typically of lower dimension than $\theta$. A particularly challenging case for uniqueness is in models with additive noise forms, such as (12); in this case it will generally be necessary to fix a prior for the variance using knowledge of the measurement process.
To handle more general cases of lack of uniqueness, we seek to encode the following principle into our methodology: given two estimates of $\lambda^{*}$ which have equivalent values of $\tilde{D}(\lambda^{*}\mid\boldsymbol{X})$, we prefer the one with the larger variance for $\text{P}(\theta\mid\lambda^{*},\boldsymbol{X})$. We accomplish this by defining a secondary objective $\tilde{N}(\lambda\mid\boldsymbol{X})$ formed from a suitable function $n(\theta)$ with
$\tilde{N}(\lambda\mid\boldsymbol{X})=\int n(\theta)\,\text{d}\text{P}(\theta\mid\lambda,\boldsymbol{X}).$ (3)
In this paper we consider only one form for this: the (negative) mean of the marginal log standard deviations of each of the $Q$ components of $\theta\in\Theta\subseteq\mathbb{R}^{Q}$.
$\tilde{N}(\lambda\mid\boldsymbol{X})=-\frac{1}{Q}\sum_{q=1}^{Q}\log\left(\text{SD}_{\,\text{P}(\theta_{q}\mid\lambda,\boldsymbol{X})}\left[\theta_{q}\right]\right),$ (4)
where $\text{SD}_{\text{P}(Z)}[Z]$ is the standard deviation of $Z$ under distribution $\text{P}(Z)$. Analytic expressions for $\text{SD}_{\text{P}(\theta\mid\lambda,\boldsymbol{X})}[\theta_{q}]$ can be used if available; or Monte Carlo estimates otherwise.
2.4 Algorithm and optimisation
We jointly minimise (1) and (4) using a multi-objective optimisation algorithm, and obtain a set of possible $\lambda$ values which comprise the Pareto frontier $\mathcal{P}=\{\lambda_{l}\}_{l=1}^{\lvert\mathcal{P}\rvert}$. This is the set of all “non-dominated” choices for $\lambda$, meaning that no point in $\mathcal{P}$ is preferable in both objectives to any of the remaining points in $\mathcal{P}$ (Deb, 2001, chap. 2). For each $\lambda$ in $\mathcal{P}$ we compute the loss
$\tilde{L}(\lambda)=\log(\tilde{D}(\lambda\mid\boldsymbol{X}))+\kappa\,\tilde{N}(\lambda\mid\boldsymbol{X}),$ (5)
where the value of $\kappa>0$ expresses our relative belief in the importance of the secondary objective. The optimal value is then $\lambda^{*}:=\min\limits_{\lambda\in\mathcal{P}}\tilde{L}(\lambda)$.
This optimum depends on $\kappa$, which will usually be difficult to assess. However, using multi-objective optimisation we can evaluate (5) for any $\kappa$ without needing to redo the optimsation step, and thus plot Pareto frontiers for a wide range of values $\kappa\in\mathcal{K}$ coloured by loss, with the minimum loss point indicated. These can guide our choice of $\kappa$: we can seek a value of $\kappa$ with the minimum loss point not on the extreme of the Pareto frontier, since we would like to balance the two objectives. This approach is particularly useful in settings where the scales of the two optima differ markedly, which we further discuss in Supplement S2. Where it is feasible to replicate the optimisation procedure, we can additionally seek a choice of $\kappa$ that leads to Pareto frontiers with minimal variability across replicates, since this suggests the optimal solution can be estimated reliably.
We use a two-stage global optimisation process. Our algorithm requires: a method for sampling $\text{P}(Y\mid\lambda,\boldsymbol{X})$; upper and lower limits that render $\Lambda$ a compact subset of $\mathbb{R}^{L}$, due to our use of global optimisation; and methods to evaluate the log-target CDF $\log(\text{T}(Y\mid\boldsymbol{X}))$ and for drawing samples according to $\text{T}(Y\mid\boldsymbol{X})$. The first optimisation stage in our algorithm considers only $\tilde{D}(\lambda\mid\boldsymbol{X})$ to focus on faithfulness, whereas the second stage also considers $\tilde{N}(\lambda\mid\boldsymbol{X})$ to account for uniqueness and replicability. We adopt this approach because minimising $\tilde{D}(\lambda\mid\boldsymbol{X})$ is considerably more challenging than minimising $\tilde{N}(\lambda\mid\boldsymbol{X})$. An idealised form of this process is illustrated in Figure 1. We briefly describe the algorithm below; full details are in Supplement S3.
Figure 1: Illustration of the algorithm, which seeks to match the prior (blue) and target distribution (red) by optimising $\lambda$. The initial value $\lambda$ produces a poor match. Stage 1 minimises (1); stage 2 then minimises (5), which increases the variance of $p(\theta\mid\lambda^{*})$.
In stage one we minimise $\tilde{D}(\lambda\mid\boldsymbol{X})$ using controlled random search 2 (CRS2) with local mutation (Kaelo & Ali, 2006), which we run for $N_{\text{CRS2}}$ iterations. We make use of the final optimum value $\lambda^{*}$, as well as the $N_{\text{CRS2}}$ trial points to obtain a design $\mathcal{D}$ for the next stage. The design comprises values of $\lambda$, and their corresponding values of $\log(\tilde{D}(\lambda\mid\boldsymbol{X}))$. A (small) number of padding points $N_{\text{pad}}$ are added to $\mathcal{D}$ for numerical robustness in stage 2. The result is the design $\mathcal{D}=\left\{\lambda_{i},\log(\tilde{D}(\lambda_{i}\mid\boldsymbol{X}))\right\}_{i=1}^{N_{\text{design}}+N_{\text{pad}}}$. Whilst CRS2 was not designed to minimise noisy functions, it appears empirically robust to small quantities of noise.
Stage one output is then used to initialise stage two, which additionally focuses on uniqueness and replicability by employing multi-objective Bayesian optimisation (Frazier, 2018) via MSPOT (Zaefferer et al., 2012) to jointly minimise $\tilde{D}(\lambda\mid\boldsymbol{X})$ and $\tilde{N}(\lambda\mid\boldsymbol{X})$. MSPOT uses a separate Gaussian process (GP) approximation to each of the objectives, and evaluates these approximations at many points from a Latin hypercube design. At each iteration the best points under the current GP approximations are evaluated using the actual objectives and used to iteratively improve the approximations. The noisy (in practice) and computationally expensive nature of our objectives, particularly $\tilde{D}(\lambda\mid\boldsymbol{X})$, necessitates an approach such as MSPOT. Employing GP models for the objectives enables inexpensive screening of values of $\lambda\in\Lambda$ that are far from optimal. Moreover, the GP is a flexible yet data efficient model to use as an approximation and can, through appropriate choice of kernel, capture correlation or other complex relationships between components of $\lambda$ and the objective. We use an optional batching technique in stage two because the computational cost of evaluating the GP grows cubically in the number of points $N_{\text{BO}}$ used in its construction. After $N_{\text{BO}}$ iterations, the evaluated points are reduced to their Pareto frontier (Kung et al., 1975). Note finding the global optimium is not guaranteed by our optimisation strategy (Mullen, 2014).
To approximate $\tilde{D}(\lambda\mid\boldsymbol{X})$ we first approximate the prior predictive CDF $\text{P}(Y\mid\lambda,\boldsymbol{X})$ by drawing $S_{r}$ samples $\boldsymbol{y}_{r}^{(\text{P})}=(y_{s,r})_{s=1}^{S_{r}}$ with $\boldsymbol{y}_{r}^{(\text{P})}\sim\text{P}(Y\mid\lambda,X_{r})$ to form the ECDF $\hat{\text{P}}(Y\mid\lambda,X_{r},\boldsymbol{y}_{r}^{(\text{P})})$, given values of $\lambda$ and $X_{r}$. We then approximate (1), denoted $D(\lambda\mid\boldsymbol{X})$, using $I_{r}$ samples $(y_{i,r})_{i=1}^{I_{r}}\sim\text{Q}(Y\mid X_{r})$ drawn from an importance distribution $\text{Q}(Y\mid X_{r})$ with
$D(\lambda\mid\boldsymbol{X})=\frac{1}{R}\sum_{r=1}^{R}\frac{1}{I_{r}}\sum_{i=1}^{I_{r}}d(\text{P}(y_{i,r}\mid\lambda,X_{r}),\text{T}(y_{i,r}\mid X_{r}))\frac{\text{d}\text{T}(y_{i,r}\mid X_{r})}{\text{d}\text{Q}(y_{i,r}\mid X_{r})}.$ (6)
We select $\text{Q}(Y\mid X_{r})$ using information about the support $\mathcal{Y}$, and samples from $\text{P}(Y\mid\lambda,X_{r})$ and $\text{T}(Y\mid X_{r})$. Approximating $\tilde{N}(\lambda\mid\boldsymbol{X})$ is usually straightforward via Monte Carlo, and we denote the corresponding estimate (or analytic form, if available) by $N(\lambda\mid\boldsymbol{X})$.
2.5 Benchmarking and other empirical considerations
We show results for both the multi-objective approach and a single-objective approach, which optimises only (1) even in Stage 2. Given $\lambda^{*}$, we empirically assess faithfulness by comparing the target distribution $\text{T}(Y\mid X_{r})$ and the estimated optimal prior predictive distribution $\text{P}(Y\mid\lambda^{*},X_{r})$. Replicability and uniqueness are more challenging to disentangle empirically: without replicability we are unable to conclude whether the multi-objective optimisation problem admits a unique solution. We will first assess replicability by examining the stability of the components of the loss in (5) across independent replications of the optimisation procedure. When the loss is stable across replicates, we will assess uniqueness by examining whether the optimal prior predictive distribution $\text{P}(Y\mid\lambda^{*},X_{r})$ and prior $\text{P}(\theta\mid\lambda^{*},\boldsymbol{X})$ are stable across replicates; stability of both is good evidence of uniqueness.
3 Examples
3.1 Calibrating a cure fraction survival model
Cure models (Amico & Van Keilegom, 2018) for survival data are useful when a cure mechanism is physically plausible a priori, and when individuals are followed up for long enough to be certain all censored individuals in our data are “cured”. Such lengthy follow ups are not always possible, but a cure model remains plausible when a large fraction of the censored observations occur after the last observed event time. However, we cannot distinguish in the right tail of the survival time distribution between censored uncured individuals and genuinely cured individuals. We suppose here that we possess prior knowledge on the fraction of individuals likely to be cured, and the distribution of event times amongst the uncured, and seek to translate this information into a prior. This setting is interesting because the target distribution is of mixed discrete/continuous type, due to censoring. Additionally, we specify a model with a nontrivial correlation structure, about which we wish to specify an informative prior, which is known to be challenging.
Target survival time distribution and covariate generation
Suppose that individuals are followed up for an average (but arbitrary) of 21 units of time, with those who experience the event doing so a long time before the end of follow up. Furthermore, suppose we believe that, a priori, $5\%$ of the patients will be cured, with $0.2\%$ of events censored due to insufficient follow up.
Consider individuals $n=1,\ldots,N$ with event times $Y_{n}$ and censoring times $C_{n}$, such that $Y_{n}\in(0,C_{n}]$. A target distribution that is consistent with our beliefs comprises a point mass of $0.05$ at $C_{n}$, and a lognormal distribution with location $\mu^{\text{LN}}=\log(3)$ and scale $\sigma^{\text{LN}}=2\mathop{/}3$ for $Y_{n}<C_{n}$. This choice of lognormal has $99.8\%$ of its mass residing below 21, and thus produces event times that are “well separated” from the censoring time, as required by a cure fraction model. Denoting the lognormal CDF with $\text{LogNormal}(Y;\mu,\sigma^{2})$, we define the target CDF
$\text{T}(Y_{n}\mid C_{n})=0.95\frac{\text{F}^{\text{LN}}(Y_{n};\mu^{\text{LN}},(\sigma^{\text{LN}})^{2})}{Z_{n}}+0.05\mathbbm{1}_{\left\{Y_{n}=C_{n}\right\}},\qquad Y_{n}\in(0,C_{n}],$ (7)
where $Z_{n}=\text{LogNormal}(C_{n};\mu^{\text{LN}},\left(\sigma^{\text{LN}}\right)^{2})$ is the required normalising constant.
We simulate data for this example with $N=50$ individuals, each with $4$ correlated covariates. When we consider the censoring time $C_{n}$, which also functions as a covariate, we have $B=5$ covariates (we use $B$ instead of $C$ as in Section 2 for clarity). In line with our target distribution, simulated censoring times are distributed such that $C_{n}\sim 20+\text{Exp}(1)$. We sample a single correlation matrix $\boldsymbol{Q}\sim\text{LKJ}(5)$ (Lewandowski et al., 2009) and subsequently covariates $\tilde{\mathbf{x}}_{n}\sim\text{MultiNormal}(\boldsymbol{0},\boldsymbol{Q})$. This results in marginally-standardised yet correlated covariates.
3.1.1 Model
A cure model for survival data, expressed in terms of its survival function, is
$S(Y\mid X,\theta)=\pi+(1-\pi)\tilde{S}(Y\mid\tilde{\boldsymbol{X}},\tilde{\theta}),$ (8)
where a proportion $\pi\in(0,1)$ of the population are cured and never experience the event of interest. The survival times for the remaining $1-\pi$ proportion of the population are distributed according to the uncured survival function $\tilde{S}(Y\mid\tilde{\boldsymbol{X}},\tilde{\theta})$. We use the tilde in $\tilde{\boldsymbol{X}}$ and $\tilde{\theta}$ to denote quantities specific to the uncured survival distribution, and denote $\theta=(\pi,\tilde{\theta})$ to align with our general notation.
Right censoring results in $Y_{n}=C_{n}$. The censoring indicator $\delta_{n}=\mathbbm{1}_{\left\{Y_{n}<C_{n}\right\}}$ is 0 for right censored events, and is 1 otherwise. We denote with $\tilde{\mathbf{x}}_{n}$ the $n$th row of the $N\times(B-1)$ covariate matrix $\tilde{\boldsymbol{X}}$, which we assume is column-wise standardised. We assume a Weibull regression model for the uncured event times, with survival function
$\begin{gathered}\tilde{S}(Y_{n}\mid\tilde{\theta},\tilde{\mathbf{x}}_{n},C_{n})=\exp\left\{-Y_{n}^{\gamma}\exp\left\{\beta_{0}+\tilde{\mathbf{x}}_{n}\boldsymbol{\beta}\right\}\right\},\quad Y_{n}\in(0,C_{n}]\end{gathered}$ (9)
with $\tilde{\theta}=(\gamma,\beta_{0},\boldsymbol{\beta})$. The likelihood, with hazard $\tilde{h}(Y_{n}\mid\tilde{\theta},\tilde{\mathbf{x}}_{n},C_{n})$, for the $n$th individual is
$\text{p}(Y_{n}\mid\theta,\tilde{\mathbf{x}}_{n},C_{n})=$ $\left((1-\pi)\tilde{S}(Y_{n}\mid\tilde{\theta},\tilde{\mathbf{x}}_{n},C_{n})\tilde{h}(Y_{n}\mid\tilde{\theta},\tilde{\mathbf{x}}_{n},C_{n})\right)^{\delta_{n}}$ (10)
×$\times\left(\pi+(1-\pi)\tilde{S}(Y_{n}\mid\tilde{\theta},\tilde{\mathbf{x}}_{n},C_{n})\right)^{1-\delta_{n}}.$
In the notation of Section 2, we have $Y=(Y_{n})_{n=1}^{N}$ and $X=(C_{n},\tilde{\mathbf{x}}_{n})_{n=1}^{N}$, with $X$ including censoring times because the support of $Y\mid X_{r}$ depends on $X_{r}$.
We will seek to identify optimal values of the hyperparamers $\lambda=(\alpha,\beta,\mu_{0},\allowbreak\sigma^{2}_{0},s_{\beta},\boldsymbol{\omega},\allowbreak\boldsymbol{\eta},\allowbreak a_{\pi},b_{\pi})^{\top}$, with $\pi\sim\text{Beta}(a_{\pi},b_{\pi})$, $\gamma\sim\text{Gamma}(\alpha,\beta)$, $\beta_{0}\sim\text{Normal}(\mu_{0},\sigma_{0}^{2})$ and $\boldsymbol{\beta}\sim\allowbreak\text{MVSkewNormal}(\boldsymbol{0},\allowbreak\boldsymbol{S},\boldsymbol{\eta})$, with $\boldsymbol{S}=\text{diag}(s_{\beta})\,\,\boldsymbol{\Omega}\,\,\text{diag}(s_{\beta})$ where $s_{\beta}$ is the prior marginal scale of $\boldsymbol{\beta}$ and $\boldsymbol{\Omega}$ is parameterised by $\boldsymbol{\omega}=(\omega_{1},\ldots,\omega_{6})^{\top}\in[-1,1]^{6}$ that uniquely determine its Cholesky factor. The skewness is necessary to incorporate the nonlinear relationship between the hazard and the effect of the covariates, and a covariance structure is used to account for fact that not all the elements of $\boldsymbol{\beta}$ can be large simultaneously. Further details are in Supplement S4.
3.1.2 Results
[…]
We use $S_{r}=10^{4}$ prior predictive samples and $I_{r}=5\times 10^{3}$ importance samples to estimate the discrepancy; and optimise using $N_{\text{CRS2}}=2000$ CRS2 iterations, followed by $N_{\text{batch}}=3$ batches of Bayesian optimisation with $N_{\text{BO}}=200$ iterations per batch, carrying $N_{\text{design}}=60$ points between batches. We select $\kappa$ by inspecting the Pareto frontiers for $\kappa\in\{0.1,0.2,0.3,0.5,1,2\}$ (Figure 2 and Supplement S4.2). Except for the maximum and minimum values, which yield minimum loss points on the extremes of the Pareto frontier, the minimum loss point is insensitive to a wide range of $\kappa$ values. We select $\kappa=0.3$ which simultaneously minimises variability in loss and both objectives.
The values of the loss and discrepancy functions at $\lambda^{*}$ across replicas are tightly distributed (see Supplement S4.3), which indicates replicability. We assess faithfulness by inspecting Figure 3. Across all replicates and discrepancies the estimated optimal prior predictive distribution is highly faithful to the target.
Figure 3: Estimated optimal prior predictive densities $\text{p}(Y_{n}\mid\lambda^{*})$ (red/blue lines and dots) and target densities $\text{t}(Y_{n}\mid C_{N})$ (black lines and crosses) for individual $n=9$ (other individuals are visually indistinguishable).
Figure 4: Estimated optimal prior marginal densities of $\text{p}(\theta\mid\lambda^{*})$ for each component of $\theta$ ($\beta_{2},\dots,\beta_{4}$, not shown, are near identical to $\beta_{1}$). Both axes are manually truncated for readability.
Figure 4 displays the marginals of $\theta$ for each independent replicate. The single objective approach (i.e. only optimising predictive discrepancy, shown in blue) consistently locates the degenerate, non-unique solution where all the variation in the uncensored event times is attributed to the baseline hazard shape $\gamma$ and the intercept $\beta_{0}$: i.e. all the mass for $\boldsymbol{\beta}$ (the regression coefficients) is close to 0. The combination of $\gamma$ and $\beta_{0}$ is evidently poorly identified, and further calculation reveals that only the derived product $\gamma\exp{\beta_{0}}$ is uniquely determined. Given the inter-replicate consistency previously observed in Supplement S4.3 we infer that the solution is not unique. In the multi-objective approach, there is a preference for the optima surrounding $\gamma\approx 7.5,\beta_{0}\approx-10$; an improvement in uniqueness over the single objective approach, but imperfect.
Figure 5 displays the bivariate prior marginal densities for $\beta_{3}$ and $\beta_{4}$, two representative elements of $\boldsymbol{\beta}$. Nonuniqueness is clearly apparent, with both positive and negative marginal skewness possible. The multi-objective approach suggests a wider distribution for $(\beta_{3},\beta_{4})$, as does the CvM discrepancy relative to the AD discrepancy.
Figure 5: Contours of the log prior density $\log(\text{p}(\beta_{3},\beta_{4}\mid\lambda^{*}))$ at the optima. For clarity we plot only the final 12 replicates, each in a unique colour.
Overall, the procedure produces priors that faithfully reflect the target, in a replicable manner. However, neither multi- or single-objective solutions are unique, particularly for the covariance structure, with the former closer to unique for $\gamma$.
3.2 Priors from model-derived quantities
Consider the linear model $Y=\boldsymbol{X}\boldsymbol{\beta}+\varepsilon$ for $n\times p$ design matrix $\boldsymbol{X}$ and $p$-vector of coefficients $\boldsymbol{\beta}$ indexed by $j=1,\ldots,p$, and where the noise $\varepsilon$ has zero mean and variance $\sigma^{2}$. Suppose information about the fraction of variance explained by the model is available – from previous similar experiments, or from knowledge of the measurement process – in the form of a plausible distribution for the coefficient of determination
$R^{2}=1-\frac{\sigma^{2}}{n^{-1}\boldsymbol{\beta}^{\top}\boldsymbol{X}^{\top}\boldsymbol{X}\boldsymbol{\beta}+\sigma^{2}}.$ (11)
assuming that the columns of $\boldsymbol{X}$ have been centred. Our aim is to use our knowledge of $R^{2}$ to set suitable priors for the regression coefficients $\beta$. This idea was the inspiration for a class of shrinkage priors (Zhang & Bondell, 2018; Zhang et al., 2022), but we would like to make this idea applicable to a wider selection of prior structures.
We consider three priors for the regression coefficients: a Gaussian prior and two shrinkage priors. To demonstrate the challenge that noise parameters pose for uniqueness we will assume $\varepsilon\sim N(0,\sigma^{2})$ and seek to select the hyperparameters for $\sigma^{2}\sim\text{InverseGamma}(a_{1},b_{1})$.
The Gaussian prior $\beta_{j}\sim\text{N}\left(0,\frac{\sigma^{2}}{\gamma}\right)$ has only one hyperparameter $\gamma$, which controls the ratio of prior variability due to $\boldsymbol{\beta}$ to that of $\varepsilon$. Hence, we denote parameters $\boldsymbol{\theta}_{\text{GA}}=(\boldsymbol{\beta},\sigma^{2})$, and seek optimum values for hyperparameters $\boldsymbol{\lambda}_{\text{GA}}=(\gamma,a_{1},b_{1})$.
The Dirichlet-Laplace prior (Dir. Lap.) is defined (Bhattacharya et al., 2015) for the $j$th coefficient such that $\beta_{j}\sim\text{Laplace}\left(0,\sigma\phi_{j}\tau\right)$, $(\phi_{1},\ldots,\phi_{p})\sim\text{Dirichlet}(\alpha,\ldots,\alpha)$, $\tau\sim\text{Gamma}(p\alpha,1\mathop{/}2)$. Smaller values of the single hyperparameter $\alpha$ yield more sparsity in $\boldsymbol{\beta}$. Thus we denote $\boldsymbol{\lambda}_{\text{DL}}=(\alpha,a_{1},b_{1})$ and $\boldsymbol{\theta}_{\text{DL}}=(\boldsymbol{\beta},\sigma^{2},\phi_{1},\ldots,\phi_{p},\tau)$.
The regularised horseshoe prior (Reg. Horse.) (Piironen & Vehtari, 2017) has more intermediary quantities and less linearity, increasing its flexiblity but making finding optimal hyperparameter values more challenging. The prior is
$\begin{gathered}c^{2}\sim\text{InvGamma}\left(\frac{\nu}{2},\frac{\nu s^{2}}{2}\right),\quad\omega\sim\text{Cauchy}^{+}\left(0,\frac{p_{0}}{p-p_{0}}\sqrt{\frac{\sigma^{2}}{n}}\right),\\ \delta_{j}\sim\text{Cauchy}^{+}(0,1),\quad\tilde{\delta}^{2}_{j}=\frac{c^{2}\delta^{2}_{j}}{c^{2}+\omega^{2}\delta^{2}_{j}},\quad\beta_{j}\sim\text{N}(0,\omega^{2}\tilde{\delta}^{2}_{j}),\end{gathered}$
with $\text{Cauchy}^{+}$ denoting a Cauchy distribution truncated to $[0,\infty)$. Whilst the regularised horseshoe is carefully designed to make $(p_{0},\nu,s^{2})$ interpretable and easy to choose, here we aim to see if we can choose $\boldsymbol{\lambda}_{\text{HS}}=(p_{0},\nu,s^{2},a_{1},b_{1})$ to match an informative prior for $R^{2}$. We denote $\boldsymbol{\theta}_{\text{HS}}=(\boldsymbol{\beta},\sigma^{2},c^{2},\omega,\delta_{1},\ldots\delta_{p})$.
3.2.1 Evaluation setup and tuning parameters
To assess each prior’s ability to faithfully encode the information present across a wide variety of target distribution and assess the uniqueness and replicability of the optimisation process, we consider sixteen different $\text{Beta}(s_{1},s_{2})$ distributions as our target $\text{T}(R^{2})$, with $\{s_{1},s_{2}\}\in\mathcal{S}\times\mathcal{S}$ and $\mathcal{S}$ chosen to be four exponentially-spaced values between and including $1\mathop{/}3$ and $3$ (i.e. equally-spaced between $\log(1\mathop{/}3)$ and $\log(3)$). These values represent a variety of potential forms of the supplied target predictive distribution for $R^{2}$.
We fix $n=50$ and $p=80$ with entries in $\boldsymbol{X}$ drawn from a standard Gaussian distribution, and assess replicability using 10 independent runs for each prior and target. The support $\Lambda$ for the hyperparameters is defined in Supplement S5.1. We use $S=10^{4}$ prior predictive samples, $I=5\times 10^{3}$ importance samples from a $\text{Uniform}(0,1)$, and use both $d^{\text{AD}}$ and $d^{\text{CvM}}$ as discrepancy functions. We employ $N_{\text{CRS2}}=1000$ iterations, and subsequently perform both single and multi-objective Bayesian optimisation for $N_{\text{batch}}=1$ batch of $N_{\text{BO}}=150$ iterations, using $N_{\text{design}}=50$ points from the first stage. The single objective approach illustrates that differences in flexibility between priors also induce differences in uniqueness, and highlights issues in choosing a prior for the additive noise parameter $\sigma^{2}$. Choosing $\kappa$ is challenging in the multi-objective approach, as its value should depend on the target, the discrepancy function, and the prior. These dependencies result in 96 possible choices of $\kappa$, which is an infeasible number of choices to make in this example. Instead we fix $\kappa=0.5$ for all multi-objective settings. We use the secondary objective (4), except for quantities where the standard deviation is undefined for some $\lambda\in\Lambda$, for which we use a robust scale estimator (Rousseeuw & Croux, 1993).
3.2.2 Results
We first assess replicability. It appears from Figure 6 that both discrepancies are replicable for the both Gaussian and Dirichlet-Laplace priors. In contrast, the results for the Regularised Horseshoe prior appear to replicate poorly under the AD discrepancy, but reasonably under the CvM discrepancy.
Figure 6: Discrepancy at the optima $\log(D(\lambda^{*}))$ for four target distributions across 10 replicates, with each the mean of 10 evaluations of $\log(D(\lambda^{*}))$ for the same $\lambda^{*}$.
We evaluate faithfulness by inspecting the densities $\text{p}(R^{2}\mid\lambda^{*})$ and $\text{t}(R^{2})$ for the various targets (all distributions in this example have corresponding densities). A selected subset of the pairs of $(s_{1},s_{2})$ values are displayed in Figure 7 (complete results are in Supplement S5.2). The Gaussian prior is universally poorly faithful. Both shrinkage priors perform better in cases where one of $s_{1}$ or $s_{2}$ is less than 1, with the regularised horseshoe performing better for the $s_{1}=s_{2}>1$ cases. Interestingly, the results are not symmetric in $s_{1}$ and $s_{2}$; the Dirichlet-Laplace prior is able to match the $s_{1}=3,s_{2}=0.69$ target well, with many of regularised horseshoe replicates performing poorly; whilst the relative performance is reversed for $s_{1}=0.69,s_{2}=3$ (see Figure 17). There is also perceptibly more variability in the regularised horseshoe replicates, which suggests the optimisation problem is more challenging and the predictive discrepancy objective is noisier. The multi-objective generally produces more variable sets of optima, which is expected, as it is a more difficult optimisation problem but we do not allow it additional computational resources. There is little visible difference between the two discrepancy functions. Finally, as the values of $s_{1}$ and $s_{2}$ increase, the faithfulness of the shrinkage priors generally decreases. Across the full set of simulations, the regularised horseshoe is evidently the most flexible.
Figure 7: Optimal prior predictive densities $\text{p}(R^{2}\mid\lambda^{*})$ for the three priors considered, for selected target densities (black lines). Density values are truncated to $[0,10]$ for readability.
To assess uniqueness, we consider estimated optimal hyperparameter values $\lambda^{*}$ in each replicate. Figure 8 displays the estimates for $s_{1}=3$ and $s_{2}\in\{0.33,0.69,1.44,3\}$, which corresponds to the targets in Figure 7. The estimates for $\gamma$ and $\alpha$, for the Gaussian and Dirichlet-Laplace priors respectively, are consistent across replicates, which suggests the optima may be unique. This remains true even for targets where the prior is not faithful to the target, e.g. the $\text{Beta}(3,3)$ target. There is more variability in the hyperparameters of the regularised horseshoe prior. There does appear to be unique solution for $\nu$ for the $\text{Beta}(3,0.33)$ and $\text{Beta}(3,0.69)$ targets, whereas $p_{0}$ and $s^{2}$ are highly variable across replicates, which may reflect nonuniqueness or may be due to the lack of replicability (discussed above) of this optimisation for the regularised horseshoe.
Figure 8: Optimal values $\lambda^{*}$ for each of the three priors considered. Columns contain (possibly prior-specific) hyperparameters, with the point colour corresponding to a specific prior. Each point’s shape corresponds to the combination of discrepancy function and single or multi-objective approach. The target beta densities (denoted by the row panel titles) correspond to Figure 7.
The hyperparameters $(a_{1},b_{1})$ for the additive noise variance $\sigma^{2}$ are highly variable across replications for almost all prior/target combinations. This reflects the anticipated lack of uniqueness when incorporating such hyperparameters. It is particularly striking for the Dirichlet-Laplace prior when $s_{2}\in\{0.33,0.69\}$, where we observe consistent and excellent fits/faithfulness to the target, but these do not correspond to replicable estimates for $(a_{1},b_{1})$. These settings are also interesting as the choice of single or multi-objective approach greatly impacts the optimum values of $a_{1}$ and $b_{1}$. Faithfulness of the multi-objective optima, illustrated in Figure 7, are not appreciably worse than the single objective approach, but the inclusion of $\sigma^{2}$ into the secondary objective has resulted in optimal values of $a_{1}$ and $b_{1}$ that maximise the dispersion of the marginal prior (i.e. small $a_{1}$ and large $b_{1}$). Asymptotic results are also known for the Gaussian prior, and in Supplement S5.3 we further assess replicability by benchmarking our optimisation process against suitable ‘true’ (asymptotically) values.
Our optimisation procedure has minimised $\log(D(\lambda))$ using both the AD and CvM discrepancy functions. The former places extra emphasis on matching the tails of the target, and thus the Regularised Horseshoe values in the top row of Figure 6 differ from our expectations given the results in the top two rows of Figure 7. Take, for example, the $s_{1}=3,s_{2}=0.69$ case. It is plainly evident from Figure 7 that the regularised horseshoe prior provides a better fit to the target distribution at $\boldsymbol{\lambda}^{*}_{\text{HS}}$, and yet the corresponding $\log(D(\boldsymbol{\lambda}_{\text{HS}}^{*}))$ values in the top row of Figure 6 suggest that it is considerably worse that the Gaussian prior at $\boldsymbol{\lambda}_{\text{GA}}^{*}$. To reconcile this apparent contradiction, we inspect $\log(D(\lambda))$ at the optima computed using the Cramér-Von Mises discrepancy function. These values are displayed in the bottom row of Figure 6, whose values closely match our expectations given Figure 7. Given the range of behaviours of $\text{p}(R^{2}\mid\lambda^{*})$ for all the optima, we can conclude that AD more heavily penalises over-estimation of the tails of $\text{p}(R^{2}\mid\lambda^{*})$ than under-estimation. This does not discount it as an optimisation objective, but does complicate comparisons between competing priors.
Overall, this example illustrates how information about a model-derived, nonobservable quantity can be used to form an informative prior. The most flexible shrinkage model (the regularised horseshoe prior) was almost always the most faithful to the supplied information. Conversely, the Gaussian prior is the most replicable and unique, but the lack of faithfulness means it is unsuitable in combination with a Beta prior on $R^{2}$.
3.3 A human-aware prior for a human growth model
Suppose an individual has their height measured at age $t_{m}$ (in years) for $m=1,\ldots,M$, with corresponding measurement $y_{m}$ (in centimetres). The first Preece-Baines model (Preece & Baines, 1978) for human height is the nonlinear regression model
$y_{m}$ $=h(t_{m};\theta)+\varepsilon_{m}$ (12)
$=h_{1}-\frac{2(h_{1}-h_{0})}{\exp\{s_{0}(t_{m}-\gamma)\}+\exp\{s_{1}(t_{m}-\gamma)\}}+\varepsilon_{m},$ (13)
with $\varepsilon_{m}\sim\text{N}(0,\sigma^{2}_{y})$. Some constraints are required to identify this model and ensure its physical plausibility: specifically, we require $0<h_{0}<h_{1}$ and $0<s_{0}<s_{1}$. To satisfy these constraints, we parameterise in terms of $\delta_{h}=h_{1}-h_{0}$ and $\delta_{s}=s_{1}-s_{0}$, which results in $(h_{0},\delta_{h},s_{0},\delta_{s})$ all sharing the same positivity constraint. We also constrain $\gamma$ such that $\gamma\in(\min_{m}(t_{m}),\max_{m}(t_{m}))$. Even with these constraints the the denominator of the fraction can be very small, yielding negative heights, meaning the model is not plausible for all parameter values. Furthermore, the model is poorly behaved under a flat prior, so prior information is required to stabilise and/or regularise the posterior.
We thus seek in this example to specify priors congruent with two specific target prior predictive distributions. We choose $\text{LogNormal}(\mu_{q},s^{2}_{q})$ priors for each of the $q=1,\ldots,5$ elements of $\theta=(h_{0},\delta_{h},s_{0},\delta_{s},\gamma)$, and seek to identify the optimal values of $\lambda=\left(\mu_{q},s^{2}_{q}\right)_{q=1}^{5}$ (see Supplement S6.1 for $\Lambda$). We fix the prior $\sigma_{y}\sim\text{LogNormal}(0,0.2^{2})$ to avoid identifiability problems (Section 3.2). We suppose our data originate from humans, uniformly distributed between ages 2 and 18, and evenly split between sexes. We first consider a covariate-independent prior predictive density $\text{t}(Y)$ with corresponding CDF $\text{T}(Y)$ for human heights across the entire age-range, derived by summarising external data. This target (Figure 10) is a mixture of 3 gamma densities specified to approximate the external data, which is multimodal due to the fact that humans grow in spurts. We also consider a covariate-specific $\text{T}(Y\mid X_{r})$, specifying Gaussian height distributions at ages $X_{r}\in(2,8,13,18)$ (see Figure 11, and Supplement S6.2 for details).
3.3.1 Comparison with Hartmann et. al. and tuning parameters
Hartmann et al. (2020) also considered this example, but elicited 6 predictive quantiles at ages $t=(0,2.5,10,17.5)$, as opposed to entire predictive distributions at ages $t=(2,8,13,18)$ as in our covariate-specific approach. We use different ages because the model is stated to be accurate for ages $\geq 2$ (Preece & Baines, 1978). Hartmann et al. (2020) also include a noise parameter in their definition of $\theta$. The exact interpretation of this parameter is complicated by their choice of Weibull likelihood, rendering the distribution of the measurement errors sensitive to conditional mean of the model (this is still the case despite their choice of Weibull parameterisation). Finally, Hartmann et al. (2020) elicit quantiles from 5 different users and report an estimated $\lambda^{*}$ for each user. These estimates (reproduced in Supplement S6.3) allow us to compare optimal the selected priors $\text{p}(\theta\mid\lambda^{*})$
For both targets we obtain $\lambda^{*}$ using both the single objective and multi-objective optimisation processes. We use only the Cramér-Von Mises discrepancy in this example, because we were not able to reliably compute the AD discrepancy due to numerical instabilities. We use $S=5\times 10^{4}$ samples from $\text{p}(Y\mid\lambda)$ and likewise $S_{r}=5\times 10^{4}$ samples from $\text{p}(Y\mid\lambda,X_{r})$ for each of the 4 values of $X_{r}$. We use $I=5\times 10^{3}$ and $I_{r}=5\times 10^{3}$ importance samples, with $N_{\text{CRS2}}=2000$ CRS2 iterations, $N_{\text{batch}}=5$ Bayesian optimisation batches each of $N_{\text{BO}}=250$ iterations, and carry forward $N_{\text{design}}=50$ points per batch. We assess replicability using 30 independent runs of each objective/target pair.
[…]
For this example, we also assess the ‘stability’ of the resulting posterior under each prior, by separately considering each of the 93 individuals in the growth data in R-package fda (Ramsay et al., 2022). We consider each individual’s data separately, rather than jointly, to heighten the importance of the prior. We measure stability by whether Stan (Stan Development Team, 2021) flags a warning, setting adapt_delta = 0.95 and max_treedepth = 12 to minimise false positives. While a lack of warnings does not imply good model behaviour, the presence of warning clearly indicates a problem. This is a form of prior sensitivity analysis, but distinct from the ideas of Roos et al. (2015) which consider only one particular realisation of the data. We include the flat, improper prior as a benchmark.
3.3.2 Results
We consider $\kappa\in\mathcal{K}=\{0.05,0.1,\ldots,0.5\}$ for the covariate-specific targets. There is notable inter-replicate variability in the Pareto frontiers for the selection of values $\kappa\in\{0.1,0.2,0.3\}$ shown in Figure 9, due to the stochastic properties of the global optimisers we employ, with some replicates totally dominated by other replicates. Following our ‘minimum variability across replicates’ heuristic (Section 2.4), we select $\kappa=0.2$ for the covariate-specific target and $\kappa=0.15$ for the covariate-independent target (See Supplement S6.4). With these values the optimisation procedure identifies optimal $\lambda^{*}$ values with reasonably, but not entirely, consistent predictive discrepancies across replicates (see Supplement S6.5).
Figure 10: The covariate-independent marginal target density $\text{t}(Y)$ (red) and prior predictive densities $\text{p}(Y\mid\lambda^{*})$ for each of the 30 replicates (blue lines).
Figure 10 displays the target and prior predictive density estimates in the covariate-independent case. The multi-objective replicates are obtained after $\kappa=0.2$ is chosen. We see that introducing the secondary objective produces estimates of $\lambda^{*}$ that are congruent with estimates from the single objective case, though the estimates are more variable. Both single and multi-objective approaches result in reasonably, but not entirely, faithful densities for $\text{p}(Y\mid\lambda^{*})$. However, most optimum priors seem to accumulates additional probability surrounding $Y=h_{1}\approx 155$, resulting in individual trajectories attaining their adult height $h_{1}$ for younger than expected ages $t$ (which we will later confirm in Figure 12).
Figure 11: Covariate-specific target densities $\text{t}(Y\mid X_{r})$ (red lines) and prior predictive densities $\text{p}(Y\mid\lambda^{*},X_{r})$ for each of the 30 replicates (blue lines).
For the covariate-specific target (Figure 11), the secondary objective introduces in some replicates outlying estimates for $\text{p}(Y\mid\lambda^{*},X_{r})$, most clearly visible for $X_{1}=2$ and $X_{2}=8$. Both the single and multi-objective approaches struggle to match the prior predictive distribution at all ages, with consistently poorer faithfulness for $X_{1}=2$. Empirically, it does not seem possible to match all four supplied target prior predictive distributions simultaneously, given the mathematical structure of the model. Lastly, because $\text{t}(Y\mid X_{1}=2)$ is substantially narrower than the other targets, it is optimal, under the Cramér-Von Mises discrepancy, to select wider priors better matching the older age target distributions.
Figure 12: Mean (solid lines) and 95% intervals (grey regions) for prior predictives $\text{p}(h(t;\theta)\mid\lambda^{*})$ for covariate-independent and covariate-specific targets for the multi- and single-objective settings; and for each replicate from Hartmann et al. (with 75% intervals). The y-axis is truncated to $(50,200)$. The red lines are 95% intervals for $\text{t}(Y\mid X_{r})$.
Both the covariate-independent and covariate-specific targets yield plausible mean growth trajectories (Figure 12). However, covariate-independent priors are highly uncertain, resulting in implausible heights having a priori support. It is interesting to note that the covariate-specific target has similar levels of uncertainty across all ages, further suggesting that the model is too inflexible to simultaneously match all the targets when they have varying variance, as here. All 5 of the priors from Hartmann et al. (2020), for a narrower uncertainty interval, are implausible in both shape and width when viewed on this scale. It also seems unlikely that these priors accurately reflect the information provided by the experts in Hartmann et al. (2020), but this information is not reported.
Figure 13: Presence/absence of Stan warnings for all individuals (columns) in the FDA package growth data and replicate prior estimates (rows). Each replicate corresponds to a run of the optimisation process and thus a different prior (except for Flat).
The different priors produce a widely ranging proportion of warning messages in Stan (Figure 13). The flat prior produces the most warnings, with some individuals particularly prone to warning messages, suggesting that their data are relatively uninformative. The Hartmann priors produce a moderate number of warnings, with some priors less prone to produce warnings (replications 1 and 5) than others for this dataset. The covariate-specific approach produces fewer warnings than the covariate-independent approach in both the single- or multi-objective cases. This reflects the additional information available in the covariate-specific setting, and that this information results in more plausible priors. Some specific replications of the covariate-independent approach produce many warnings, suggesting these priors are inappropriate for many individuals.
Figure 14: Priors (blue) for $h_{0}$ with covariate-independent and covariate-specific targets; a flat prior scenario (prior not shown); and Hartmann et al. (2020). The corresponding posteriors for each prior for individual $n=26$ is shown in red.
The priors for $h_{0}$ exhibit substantial variability across replicates (Figure 14; see Supplement S6.6 for all $\theta$). There appear to be two distinct unimodal priors for $h_{0}$ with similar loss, suggesting that $\text{T}(Y\mid\boldsymbol{X})$ does not provided enough information to uniquely determine a prior distribution. However both priors are significantly broader than the Hartmann et. al. priors. Figure 14 also shows the posteriors for these parameters when using the (uninformative and thus challenging) data from individual $n=26$. The flat prior produces a multimodal posterior for our parameters of interest, demonstrating the lack of stability when computing the posterior and the need for prior information. As a result, the posteriors are under all methods strongly depend on the prior distribution used, which as noted is not stable under any method here. However both of our priors produce posteriors for $\delta_{s}$ with no significant mass above 2 (Supplement S6.6), as is desirable, because this corresponds physiologically implausible growth spurts that are unsupported by the data.
In summary, the priors estimated by our procedure in this example are broadly faithful to the supplied information, except in the covariate-specific case where model inflexibility prevents matching both $t=2$ and $t=18$ targets simultaneously. The covariate-specific, multi-objective method appears the most useful prior, but is arguably over concentrated, which occasionally prevents the model from fitting the data well, although all our priors successfully regularise the posterior sufficiently to enable accurate posterior sampling. Our approach does not produce a unique prior, although the secondary objective leads to a small improvement in uniqueness (see Supplement S6.6). However, some of this non-uniqueness may be attributable to imperfect replicability of the optimisation.
4 Conclusion
Setting priors for models congruent with our knowledge is often difficult without a method for translation such as we have proposed. The Preece-Baines model is a typical example, in which the observable is well understood but the model parameters are not. Similarly we anticipate our approach will be valuable for model-derived quantities (such as $R^{2}$), which are often readily reasoned about but difficult to set priors for.
One limitation of the current work is that we only partly address non-uniqueness, but we emphasise that our methodology remains valuable in such settings. Specifically, our approach provides insight into consequences of certain $\text{T}(Y\mid\boldsymbol{X})$: it facilitates identifying which components of $\lambda$ are well identified, and consideration of whether any difference between $\text{T}(Y\mid X_{r})$ and $\text{P}(Y\mid\lambda^{*},X_{r})$ are attributable to model inflexibility or an implausible target. We also have the opportunity to re-assess whether we have information that we could employ to fix certain components within $\lambda$ (e.g. the fixed prior for the noise in the human height example). Another limitation of our current work is that global optimisation methods lack guarantees of finding the global optimum in finite time, so the generalisability of the performance of our optimisation process requires further investigation in the future.
Acknowledgments and data availability
We thank Daniela De Angelis and Mevin Hooten for their feedback; and the The Alan Turing Institute under the UK Engineering and Physical Sciences Research Council (EPSRC) [EP/N510129/1] and the UK Medical Research Council [MC_UU_00002/2 and MC_UU_00002/20] for support. No original data were generated; the fda package for R contains the growth data. R code for the examples is at https://gitlab.com/andrew-manderson/pbbo-paper. For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.
S1 R package
We implement our methodology in an R package (R Core Team, 2023) pbbo (https://github.com/hhau/pbbo). pbbo builds on top of mlrMBO (Bischl et al., 2018) for multi-objective Bayesian optimisation, nlopt and nloptr (Ypma et al., 2022; Johnson, 2014) for global optimisation using CRS2 (Kaelo & Ali, 2006), and other packages for internal functionality and logging (Wickham et al., 2019; Rowe, 2016; Maechler et al., 2021). The code to reproduce the examples is available at https://gitlab.com/andrew-manderson/pbbo-paper.
S2 Further notes on choosing $\kappa$
Advantages of multi-objective optimisation are most immediately apparent when the scales of our objectives differ markedly. Consider the equivalent linearised approach, where we select $\kappa$ before optimisation and directly optimise $\tilde{L}(\lambda\mid\boldsymbol{X})$. It is generally not possible to know the range of the values of $\tilde{D}(\lambda\mid\boldsymbol{X})$ and $\tilde{N}(\lambda\mid\boldsymbol{X})$ before optimisation. Selecting an appropriate $\kappa$ without this knowledge is prohibitively difficult, leaving only the computationally expensive trial-and-error approach – where we re-run the optimiser for each new possible value of $\kappa$ – as a plausible strategy for choosing $\kappa$. In contrast, given $\mathcal{P}$ it is computationally trivial to recompute $\lambda^{*}$ for many possible values of $\kappa$ after optimisation (e.g. each panel of Figure 9 in the main text is trivial to compute). We can thus select $\kappa$ in a problem-specific manner for practically no additional computational cost to that of the multi-objective optimiser. Note that the multi-objective optimisation approach is more expensive that the linearised approach, but this additional cost is dwarfed by the number of re-runs of the latter typically required to select $\kappa$.
S3 Algorithm and optimisation details
Here we provide further details on the algorithm and optimisation process used, before providing an overarching algorithm for the complete methodology (Section S3.5)
S3.1 CRS2 as an initialiser for Bayesian optimisation
Algorithm 1 describes our use of CRS2 (Kaelo & Ali, 2006) to obtain a suitable design to initialise the Bayesian multi-objective optimisation approach in step 2.
Algorithm 1 Using CRS2 to find an initial design for Bayesian optimisation
1Log total predictive discrepancy $\log(D(\lambda\mid\boldsymbol{X}))$ (evaluable using Algorithm 4), number of CRS2 iterations to run $N_{\text{CRS2}}$, number of points in final design $N_{\text{design}}$, number of additional padding points to add for numerical stability $N_{\text{pad}}$, hyperparameter support $\Lambda$
2
3function Initial design($N_{\text{CRS2}},N_{\text{design}},N_{\text{pad}}$)
4     Initialise $\mathcal{S}=\{\}$, an empty set to hold possible design points
5     for $i$ in $1\ldots N_{\text{CRS2}}$ do
6         Minimising $\log(D(\lambda\mid\boldsymbol{X}))$, get the $i$th trial point $\tilde{\lambda}_{i}$ and value $\log(D(\lambda_{i}\mid\boldsymbol{X}))$ from CRS2 with local mutation (Kaelo & Ali, 2006)
7         Compute $\tilde{w}_{i}=-\exp\left\{\log\left(D(\tilde{\lambda}_{i}\mid\boldsymbol{X})\right)\right\}$
8         Concatenate $\mathcal{S}=\mathcal{S}\cup\left\{\tilde{\lambda}_{i},\log(D(\tilde{\lambda}_{i}\mid\boldsymbol{X})),\tilde{w}_{i}\right\}$
9     end for
10     Normalise weights such that $w_{i}=\exp\left\{\tilde{w}_{i}-\log\left(\sum_{i=1}^{N_{\text{CRS2}}}\exp\left\{\tilde{w}_{i}\right\}\right)\right\}$
11     Subsample without replacement $N_{\text{design}}$ values from $\mathcal{S}$ according to the normalised weights, and store in $\mathcal{D}=\{\lambda_{i},\log(D(\lambda_{i}\mid\boldsymbol{X}))\}_{i=1}^{N_{\text{design}}}$
12     Sample $N_{\text{pad}}$ points from a Latin hypercube design spanning $\Lambda$ (Stein, 1987), evaluate $\log(D(\lambda\mid\boldsymbol{X}))$ at these points, and add them to $\mathcal{D}$
13     return: $\mathcal{D}=\{\lambda_{i},\log(D(\lambda_{i}\mid\boldsymbol{X}))\}_{i=1}^{N_{\text{design}}+N_{\text{pad}}}$
14end function
S3.2 MSPOT
Algorithm 2 describes, in our notation, the MSPOT (Zaefferer et al., 2012) algorithm for two objectives. Note that within the algorithm we suppress each objective’s dependence on $\boldsymbol{X}$ for brevity.
Algorithm 2 Global two-objective Bayesian optimisation using MSPOT (Zaefferer et al., 2012)
1Primary objective $D(\lambda)$, secondary objective $N(\lambda)$, initial design $\mathcal{D}=\left\{\lambda_{i},D(\lambda_{i}),N(\lambda_{i})\right\}_{i=1}^{N_{\text{design}}+N_{\text{pad}}}$, number of iterations $N_{\text{BO}}$, number of new points to evaluate the surrogate models at $N_{\text{new}}$, number of evaluations to add to the design within an iteration $N_{\text{eval}}$, hyperparameter support $\Lambda$
2
3function Bayesian optimisation using MSPOT($N_{\text{BO}}$)
4     for $i$ in $1\ldots N_{\text{BO}}$ do
5         Form Gaussian process (GP) approximations to $D(\lambda)$ and $N(\lambda)$ using $\mathcal{D}$
6         Generate a new Latin hypercube design $\mathcal{N}$ of size $N_{\text{new}}$ covering $\Lambda$, such that $N_{\text{new}}\gg N_{\text{design}}$
7         for $k$ in $1\ldots N_{\text{new}}$ do
8              Use the GPs to estimate $\hat{D}(\lambda_{k})$ and $\hat{N}(\lambda_{k})$
9              Add these to $\mathcal{N}$ so that $\mathcal{N}_{k}=\left\{\lambda_{k},\hat{D}(\lambda_{k}),\hat{N}(\lambda_{k})\right\}$
10         end for
11         Truncate $\mathcal{N}$ to $N_{\text{eval}}$ points according to the non-dominated sorting rank and hypervolume contribution (Beume et al., 2007; Deb, 2001; Deb et al., 2002; Beume et al., 2009) of each point in $\left\{D(\lambda_{k}),N(\lambda_{k})\right\}_{k=1}^{N_{\text{new}}}$ with $N_{\text{eval}}\ll N_{\text{new}}$
12         for $j$ in $1\ldots N_{\text{eval}}$ do
13              Evaluate the objectives $D(\lambda_{j})$ and $N(\lambda_{j})$ for $\lambda_{j}\in\mathcal{N}$
14              Add these evaluations to $\mathcal{D}=\mathcal{D}\cup\left\{\lambda_{j},D(\lambda_{j}),N(\lambda_{j})\right\}$
15         end for
16     end for
17     Compute the Pareto frontier $\mathcal{P}=\left\{\lambda_{i},D(\lambda_{i}),N(\lambda_{i})\right\}_{i=1}^{\lvert\mathcal{P}\rvert}$ from $\mathcal{D}=\left\{\lambda_{i},D(\lambda_{i}),N(\lambda_{i})\right\}_{i=1}^{N_{\text{design}}+N_{\text{pad}}+N_{\text{BO}}N_{\text{eval}}}$ (Kung et al., 1975, see)
18     return: $\mathcal{P}$ and $\mathcal{D}$
19end function
S3.3 Inter batch resampling
Algorithm 3 describes our inter-batch resampling algorithm that we occasionally adopt in stage two of our optimisation process.
Algorithm 3 Resample the outputs from a previous batch to obtain a design for the current one.
1Pareto frontier $\mathcal{P}=\left\{\lambda_{i},\log(D(\lambda_{i}\mid\boldsymbol{X})),N(\lambda_{i}\mid\boldsymbol{X})\right\}_{i=1}^{\lvert\mathcal{P}\rvert}$ and all evaluated points =$\mathcal{E}=\left\{\lambda_{i},\log(D(\lambda_{i}\mid\boldsymbol{X})),N(\lambda_{i}\mid\boldsymbol{X})\right\}_{i=1}^{\lvert\mathcal{E}\rvert}$ from previous batch (with $\lvert\mathcal{P}\rvert\ll\lvert\mathcal{E}\rvert$), number of design points $N_{\text{design}}$, number of padding points $N_{\text{pad}}$, hyperparameter support $\Lambda$
2
3function Next batch design($N_{\text{design}},N_{\text{pad}}$)
4     Initialise $\mathcal{D}=\mathcal{P}$
5     Compute the weights $w_{i}$ for all points in $\mathcal{E}$ in the same manner as Algorithm 1 so that =$\mathcal{E}=\left\{\lambda_{i},\log(D(\lambda_{i}\mid\boldsymbol{X})),N(\lambda_{i}\mid\boldsymbol{X}),w_{i}\right\}_{i=1}^{\lvert\mathcal{E}\rvert}$
6     Sample without replacement $\max\left(N_{\text{design}}-\lvert\mathcal{P}\rvert,0\right)$ points from $\mathcal{E}$ according to the weights and add these points to $\mathcal{D}$
7     Sample $N_{\text{pad}}$ points from a Latin hypercube design covering $\Lambda$ and add these to $\mathcal{D}$
8     return: $\mathcal{D}$ such that $\lvert\mathcal{D}\rvert=\max(N_{\text{design}},\lvert\mathcal{P}\rvert)+N_{\text{pad}}$
9end function
S3.4 Evaluating $D(\lambda\mid\boldsymbol{X})$
Algorithm 4 summarises the algorithm used to evaluate $D(\lambda\mid\boldsymbol{X})$, with further explanation in the following subsections.
Algorithm 4 Evaluating approximate log total predictive discrepancy $\log(D(\lambda\mid\boldsymbol{X}))$
1Targets $\text{T}(Y\mid X_{r})$ for $r=1,\ldots,R$; samplers for generating points from $\text{T}(Y\mid X_{r})$ and $\text{P}(Y\mid\lambda,X_{r})$; discrepancy $d(\cdot,\cdot)$; number of samples to draw $S_{r}$; number of importance samples $I_{r}$; observable support $\mathcal{Y}$
2
3function Evaluate log $(D$($\lambda\mid\boldsymbol{X})$)
4     for $r$ in $1\ldots R$ do
5         Sample prior predictive $\boldsymbol{y}^{(\text{P})}_{r}=(y_{s,r}^{(\text{P})})_{s=1}^{S_{r}}\sim\text{P}(Y\mid\lambda,X_{r})$
6         Use $\boldsymbol{y}^{(\text{P})}_{r}$ to form the ECDF $\hat{\text{P}}(Y\mid\lambda,X_{r},\boldsymbol{y}^{(\text{P})}_{r})$
7         Sample target $\boldsymbol{y}^{(\text{T})}_{r}=(y_{s,r}^{(\text{T})})_{s=1}^{S_{r}}\sim\text{T}(Y\mid X_{r})$
8         Choose importance distribution $\text{Q}(Y\mid X_{r})$ via Supplement S3.4.1
9         Sample importance points $(y_{i,r})_{i=1}^{I_{r}}\sim\text{Q}(Y\mid X_{r})$
10     end for
11     Compute $\log(D(\lambda\mid\boldsymbol{X}))$ using Equations (14) – (17) in Supplement S3.4.2
12     return: Value of $\log(D(\lambda\mid\boldsymbol{X}))$
13end function
S3.4.1 Choosing importance distribution Q
Appropriate importance distributions are crucial to obtaining an accurate and low variance estimate of $D(\lambda\mid\boldsymbol{X})$. For values of $\lambda$ far from optimal, $\text{P}(Y\mid\lambda,\boldsymbol{X})$ can differ considerably from $\text{T}(Y\mid\boldsymbol{X})$. Given a specific $X_{r}$ we require an importance distribution $\text{Q}(Y\mid X_{r})$ that places substantial mass in the high probability regions of both $\text{T}(Y\mid X_{r})$ and $\text{P}(Y\mid\lambda,X_{r})$, as it is in these regions that $d(\cdot,\cdot)$ is largest. But we cannot exert too much effort on finding these densities as they are specific to each value of $\lambda$, and must be found anew for each $\lambda$.
We use three quantities to guide our choice of $\text{Q}(Y\mid X_{r})$, these being the support $\mathcal{Y}$, the samples $\boldsymbol{y}_{r}^{(\text{P})}\sim\text{P}(Y\mid\lambda,X_{r})$, and the samples $\boldsymbol{y}_{r}^{(\text{T})}\sim\text{T}(Y\mid X_{r})$. Of primary concern is the support. If $\mathcal{Y}=\mathbb{R}$ then we use a mixture of Student-$t_{5}$ distributions; for $\mathcal{Y}=\mathbb{R}=(0,\infty)$ we employ a mixture of gamma distributions; and for $\mathcal{Y}=(0,a]$ with known $a$, we opt for a mixture of Beta distributions with a discrete component at $Y=a$. The parameters of the mixture components are estimated using the method of moments. Specifically, denoting the empirical mean of $\boldsymbol{y}_{r}^{(\text{P})}$ as $\hat{\mu}^{(\text{P})}$ and the empirical variance by $\hat{v}^{(\text{P})}$, with $\hat{\mu}^{(\text{T})}$ and $\hat{v}^{(\text{T})}$ defined correspondingly for $\boldsymbol{y}_{r}^{(\text{T})}$, Table 1 details our method of moments estimators for the mixture components.
In this paper we limit ourselves to one dimensional $\mathcal{Y}$, where importance sampling is mostly well behaved or can be tamed using a reasonable amount of computation. This covers many models, and with the covariate-specific target it includes regression models. It is harder to elicit $\text{T}(Y\mid\boldsymbol{X})$ for higher dimensional data spaces, and the difficulties with higher dimensional importance sampling are well known.
$\mathcal{Y}$ $\text{Q}_{r}(Y)$ Parameter estimates Mixture weights
Notes
$\mathbb{R}$ $&\pi_{1}\text{Student-}t_{5}(Y;\hat{\mu}_{1},\hat{s}_{1})+\\ &\pi_{2}\text{Student-}t_{5}(Y;\hat{\mu}_{2},\hat{s}_{2})\end{aligned}$ $\hat{\mu}_{1}=\hat{\mu}^{(\text{P})},&\,\hat{s}_{1}=c\sqrt{\hat{v}^{(\text{P})}}\\ \hat{\mu}_{2}=\hat{\mu}^{(\text{T})},&\,\hat{s}_{2}=c\sqrt{\hat{v}^{(\text{T})}}\\ \end{aligned}$ $\pi_{1}=\pi_{2}=0.5$
$c$ defaults to $1.05$
$(0,\infty)$ $&\pi_{1}\text{Gamma}(Y;\hat{\alpha}_{1},\hat{\beta}_{1})+\\ &\pi_{2}\text{Gamma}(Y;\hat{\alpha}_{2},\hat{\beta}_{2})\end{aligned}$ $\hat{\alpha}_{1}=\frac{(\hat{\mu}^{(\text{P})})^{2}}{\tilde{\omega}^{(\text{P})}},&\,\hat{\beta}_{1}=\frac{\hat{\mu}^{(\text{P})}}{\tilde{\omega}^{(\text{P})}}\\ \hat{\alpha}_{2}=\frac{(\hat{\mu}^{(\text{T})})^{2}}{\tilde{\omega}^{(\text{T})}},&\,\hat{\beta}_{2}=\frac{\hat{\mu}^{(\text{T})}}{\tilde{\omega}^{(\text{T})}}\\ \end{aligned}$ $\pi_{1}=\pi_{2}=0.5$
$\tilde{\omega}=\min(c^{2}\hat{v},10^{5})$,
$c$ defaults to $1.05$
$[0,a]$ $&\frac{\pi_{1}}{a}\text{Beta}\left(\frac{Y}{a};\hat{a}_{1},\hat{b}_{1}\right)+\\ &\frac{\pi_{2}}{a}\text{Beta}\left(\frac{Y}{a};\hat{a}_{2},\hat{b}_{2}\right)+\\ &\pi_{3}\mathbbm{1}_{\{Y=a\}}\end{aligned}$ $\hat{a}_{1}&=\hat{\mu}^{(\text{P})}\left[\frac{\hat{\mu}^{(\text{P})}}{\tilde{\omega}^{(\text{P})}}(1-\hat{\mu}^{(\text{P})})-1\right]\\ \hat{b}_{1}&=\frac{(1-\hat{\mu}^{(\text{P})})}{\hat{\mu}^{(\text{P})}}\hat{a}_{1}\\ \hat{a}_{2}&=\hat{\mu}^{(\text{T})}\left[\frac{\hat{\mu}^{(\text{T})}}{\tilde{\omega}^{(\text{T})}}(1-\hat{\mu}^{(\text{T})})-1\right]\\ \hat{b}_{2}&=\frac{(1-\hat{\mu}^{(\text{T})})}{\hat{\mu}^{(\text{T})}}\hat{a}_{2}\end{aligned}$ $\begin{gathered}\pi_{1}=\pi_{2}=0.45\\ \pi_{3}=0.05\end{gathered}$
$\tilde{\omega}=\max(c^{2}\hat{v},10^{-6})$,
$c$ defaults to $1.05$
Table 1: Importance distributions and method of moments estimators for their constituent parametric distributions. Note that $c$ is a user-selected tuning parameter to enable the construction of wider importance distributions.
S3.4.2 Numerical considerations
For both numerical stability and optimisation performance (Eriksson & Poloczek, 2021; Snoek et al., 2014) we evaluate $D(\lambda\mid\boldsymbol{X})$ on the log scale. This is because far from optimal values of $\lambda$ have corresponding $D(\lambda\mid\boldsymbol{X})$ many orders of magnitude larger than near optimal values of $\lambda$. Furthermore, the Gaussian process approximation that underlies Bayesian optimisation assumes constant variance, necessitating a log or log-like transformation.
Suppose again that we sample $\boldsymbol{y}_{r}^{(\text{P})}\sim\text{P}(Y\mid\lambda,X_{r})$, from which we form the ECDF $\hat{\text{P}}(Y\mid\lambda,X_{r},\boldsymbol{y}_{r}^{(\text{P})})$. Having selected an appropriate importance distribution $\text{Q}(Y\mid X_{r})$ and density $\text{q}(Y\mid X_{r})$ using Supplement S3.4.1, and sample importance points $(y_{i,r})_{i=1}^{I_{r}}\sim\text{Q}(Y\mid X_{r})$, we define the intermediary quantity $z(y_{i,r})$ (in the case when densities for the target and important distribution exist, to avoid notational complexity) as
$z(y_{i,r})=\log\left(d\left(\hat{\text{P}}(y_{i,r}\mid\lambda,X_{r},\boldsymbol{y}_{r}^{(\text{P})}),\text{T}(y_{i,r}\mid X_{r})\right)\right)+\log\left(\text{t}(y_{i,r}\mid X_{r})\right)-\log\left(\text{q}(y_{i,r}\mid X_{r})\right),$ (14)
and then rewrite (6) in the main text to read
$\log(D(\lambda\mid\boldsymbol{X}))=-\log(R)+\log\left(\sum_{r=1}^{R}\exp\left\{-\log(I_{r})+\log\left(\sum_{i=1}^{I_{r}}\exp\left\{z(y_{i,r})\right\}\right)\right\}\right).$ (15)
All $\log(\sum\exp\{\cdot\})$ terms are computed using the numerically stable form (Blanchard et al., 2021).
Accurately evaluating $\log(d(\cdot,\cdot))$ in (14) involves managing the discrete nature of the ECDF (that it returns exactly zero or one for some inputs), and using specialised functions for each discrepancy to avoid issues with floating point arithmetic. We compute $\log(d^{\text{CvM}}(\cdot,\cdot))$ using
$\log\left(d^{\text{CvM}}\left(\hat{\text{P}}(y_{i,r}\mid\lambda,X_{r},\boldsymbol{y}_{r}^{(\text{P})}),\text{T}(y_{i,r}\mid X_{r})\right)\right)=2\log\left(\left\lvert\hat{\text{P}}(y_{i,r}\mid\lambda,X_{r},\boldsymbol{y}_{r}^{(\text{P})})-\exp\{\mathcal{T}(y_{i,r}\mid X_{r})\}\right\rvert\right),$ (16)
where $\mathcal{T}(y_{i,r}\mid X_{r})=\log(\text{T}(y_{i,r}\mid X_{r}))$. The log-CDF (LCDF) is often more numerically accurate for improbable values of $y_{i,r}$, and so our methodology assumes that it is this LCDF form in which the target distribution is supplied. However, because the ECDF can return exact zero/one values there is no way to perform this computation on the log scale. We thus employ high precision floating point numbers when exponentiating the LCDF values, using Rmpfr (Maechler et al., 2021), to avoid evaluating $\log(0)$.
For $\log(d^{\text{AD}}(\cdot,\cdot))$, additional care must be taken as the denominator of $d^{\text{AD}}$ in (2) in the main text tends to underflow to zero. Thus we evaluate it using
$\end{multlined}\log\left(d^{\text{AD}}\left(\hat{\text{P}}(y_{i,r}\mid\lambda,X_{r},\boldsymbol{y}_{r}^{(\text{P})}),\text{T}(y_{i,r}\mid X_{r})\right)\right)=\\ 2\log\left(\left\lvert\hat{\text{P}}(y_{i,r}\mid\lambda,X_{r},\boldsymbol{y}_{r}^{(\text{P})})-\exp\{\mathcal{T}(y_{i,r}\mid X_{r})\}\right\rvert\right)-\mathcal{T}(y_{i,r}\mid X_{r})-\texttt{log1mexp}(-\mathcal{T}(y_{i,r})),$ (17)
[…]
S3.5 Summary of complete methodology
Lastly, Algorithm 5 summaries the entire methodology we introduce in this paper.
Algorithm 5 Methodology to translate prior predictive information into a prior for the parameters in a complex model
[…]
2
3function pbbo($\kappa,N_{\text{BO}},N_{\text{batch}}$)
4     Minimising $\log(D(\lambda\mid\boldsymbol{X}))$ alone, compute the initial design $\mathcal{D}$ using CRS2 via Algorithm 1
5     for $b$ in $1\ldots N_{\text{batch}}$ do
6         Jointly minimising $\log(D(\lambda\mid\boldsymbol{X}))$ and $N(\lambda\mid\boldsymbol{X})$, compute the $b$th Pareto Frontier $\mathcal{P}_{b}$ and complete design $\mathcal{D}_{b}$ using Algorithm 2, initialising with design $\mathcal{D}$
7         Update design $\mathcal{D}$ using $\mathcal{P}_{b}$ and $\mathcal{D}_{b}$ via Algorithm 3
8     end for
9     With final Pareto frontier $\mathcal{P}_{N_{\text{batch}}}$, compute $\lambda^{*}=\min L(\lambda)=\min\limits_{\lambda\in\mathcal{P}_{N_{\text{batch}}}}\log(D(\lambda\mid\boldsymbol{X}))+\kappa N(\lambda\mid\boldsymbol{X})$
10     return: $\lambda^{*}$
11end function
S4 Additional information for the cure fraction survival example
Note that the standardisation of $\tilde{\boldsymbol{X}}$ allows us to use only one $s_{\beta}$ instead of one per covariate. These elements are transformed into $\boldsymbol{\Omega}$ using the partial correlation method of Lewandowski et al. (2009), also employed by the Stan math library (Stan Development Team, 2022). The $(B-1)$-vector $\boldsymbol{\eta}$ controls, but is not equal to, the marginal skewness for each element of $\boldsymbol{\beta}$ using the multivariate skew-normal definition of Azzalini & Valle (1996), as implemented in the sn package (Azzalini, 2022).
S4.1 Hyperparameter support $\Lambda$
See Table 2
Hyperparameter Lower Upper # Elements
$\alpha$ $\epsilon$ 20 1
$\beta$ $\epsilon$ 20 1
$\mu_{0}$10 10 1
$\sigma_{0}$ $\epsilon$ 10 1
$s_{\beta}$ $\epsilon$ 10 1
$\boldsymbol{\omega}$1 + $\epsilon$ 1 - $\epsilon$ 6
$\boldsymbol{\eta}$5 5 4
$a_{\pi}$ 1 50 1
$b_{\pi}$ 1 50 1
Table 2: Hyperparameters $\lambda$ for the cure fraction model, their upper and lower limits that define $\Lambda$, and the number of elements in the hyperparameter (which is 1 for all scalar quantities). Note that $\epsilon=10^{-4}$ is added or subtracted to the limits to avoid degenerate estimates for $\lambda$.
S4.2 Pareto Frontiers for Cramér-von Mises discrepancy
See Figure 15.
[…]
S4.3 Final objective values
See Figure 16.
Figure 16: Estimates of $D(\lambda^{*}\mid\boldsymbol{X})$, $N(\lambda^{*}\mid\boldsymbol{X})$ and $L(\lambda^{*}\mid\boldsymbol{X})$ across replicates for the cure fraction survival model.
S5 Additional information for the $R^{2}$ example
S5.1 Hyperparameter support $\Lambda$ – faithfulness experiment
See Table 3. Note that for the Dirichlet-Laplace prior, Zhang & Bondell (2018) suggest bounding $\alpha\in[(\max(n,p))^{-1},1\mathop{/}2]$. In our experiments we regularly encountered optimal values of $\alpha$ on the lower boundary, so we use instead $1\mathop{/}(3\max(n,p))$ as a lower bound.
Prior Hyperparameter Lower Upper
Gaussian $a_{1}$ 2 500
Gaussian $b_{1}$ 0.2 500
Gaussian $\gamma$ 1 500
Dir. Lap. $a_{1}$ 2 500
Dir. Lap. $b_{1}$ 0.2 500
Dir. Lap. $\alpha$ $1\mathop{/}(3\max(n,p))$ $1\mathop{/}2$
Reg. Horse. $a_{1}$ 2 500
Reg. Horse. $b_{1}$ 0.2 500
Reg. Horse. $p_{0}$ 1 $p\mathop{/}2$
Reg. Horse. $\nu$ 1 80
Reg. Horse. $s^{2}$ $10^{-5}$ 100
Table 3: Hyperparameters $\lambda$ for the $R^{2}$ example and their upper/lower limits that define $\Lambda$.
S5.2 Full faithfulness results
The complete results from the faithfulness experiment are displayed in Figure 17.
Figure 17: As in Figure 7 but for all values of $(s_{1},s_{2})$ denoted in the facet panels titles. The performance of the regularised horseshoe is superior to the Dirichlet-Laplace, both of which are vast improvements over the Gaussian.
S5.3 A comparison to an asymptotic result
The poor fit for the Gaussian prior observed in Figure 7 in the main text could be attributed to issues in the optimisation process, or to the lack of flexibility in the prior. To investigate, we compare the results for $\lambda_{\text{GA}}$ to Theorem 5 of Zhang & Bondell (2018), which is an asymptotic result regarding the optimal value of $\lambda_{GA}$ for a target $\text{Beta}(s_{1},s_{2})$ density for $R^{2}$. We compare pairs of $(n_{k},p_{k})$ for $k=1,\ldots,5$, noting that assumption (A4) of Zhang and Bondell requires that $p_{k}=\text{o}(n_{k})$ as $k\rightarrow\infty$ (for strictly increasing sequences $p_{k}$ and $n_{k}$). Thus we consider values of $p$ such that $p_{1}=80$ with $p_{k}=2p_{k-1}$ and $n$ with $n_{1}=50$ and $n_{k}=n_{k-1}^{1.2}$, both for $k=2,\ldots,5$. Each $(n_{k},p_{k})$ pair is replicated 20 times, and for each replicate we generate a different $\boldsymbol{X}$ matrix with standard normal entries. As the target density we choose $s_{1}=5,s_{2}=10$ – a “more Gaussian” target than previously considered and thus, we speculate, possibly more amenable to translation with a Gaussian prior for $\beta$. We also use this example as an opportunity to assess if there are notable differences between the Cramér-Von Mises discrepancy and the Anderson-Darling discrepancy as defined in (2) in the main text. The support $\Lambda$ for $\lambda_{\text{GA}}$ differs slightly from the example in the main text, and is defined in Table 4, as matching our target with larger design matrices requires considerably larger values of $\gamma$.
The computation of $R^{2}$ becomes increasingly expensive as $n_{k}$ and $p_{k}$ increase, which limits the value of some of our method’s tuning parameters. The approximate discrepancy function uses $S=2000$ samples from the prior predictive and is evaluated using $I=500$ importance samples. We run CRS2 for $N_{\text{CRS2}}=500$ iterations, using $N_{\text{design}}=50$ in the initial design for the subsequent single batch of Bayesian optimisation, which uses $N_{\text{BO}}=100$ iterations.
Prior Hyperparameter Lower Upper
Gaussian $a_{1}$ $2+10^{-6}$ 50
Gaussian $b_{1}$ 0.2 50
Gaussian $\gamma$ 1 5000
Table 4: Hyperparameters $\lambda$ for the asymptotic example, and their upper/lower limits that define $\Lambda$.
S5.3.1 Results and analytic comparison
Figure 18 displays the results in terms of the normalised difference between the $\gamma$ we estimate $\gamma_{\text{pbbo}}^{*}$, and the asymptotic result of Zhang and Bondell $\gamma_{\text{asym}}^{*}$. Our typical finite sample estimate is slightly larger than the asymptotic result, and the difference increases with $n_{k}$ and $p_{k}$. The variability of the normalised difference remains roughly constant, and thus reduces on an absolute scale, though extrema seem to occur more frequently for larger $n_{k}$ and $p_{k}$. These simulations suggest that the asymptotic regime has not been reached even at the largest $n_{k}$ and $p_{k}$ values we assessed.
Figure 18: Relative difference between the value of $\gamma$ obtained using our methodology ($\gamma_{\text{pbbo}}^{*}$) and Theorem 5 of Zhang and Bondell (2018) ($\gamma_{\text{asym}}^{*}$).
The estimates of $\gamma$ are not themselves particularly illuminating: we should instead look for differences in the distribution of $R^{2}$ at the optima, which is to say on the “data” scale. Figure 19 displays the target distribution and the prior predictive distribution at the optima $\text{p}(R^{2}\mid\lambda^{*}_{GA})$. The fit is increasingly poor as $n$ and $p$ increase, and there is little difference both between the two discrepancies and with each discrepancies replications. The lack of difference implies that the optimisation process is consistently locating the same minima for $D(\lambda)$. We conclude that either 1) the ability of the model to match the target depends on there being additional structure in $\boldsymbol{X}$, or 2) it is not possible to encode the information in a $\text{Beta}(5,10)$ prior for $R^{2}$ into the Gaussian prior.
Figure 19: The target density $\text{t}(R^{2})$ and optimal prior predictive densities $\text{p}(R^{2}\mid\lambda^{*})$ under both the Cramér-von Mises (red, left column) and Anderson-Darling (blue, right column) discrepancies. There are 20 replicates of each discrepancy in this plot.
This example also further illustrates the difficulties inherent in acquiring a prior for additive noise terms. Specifically, in this example it is difficult to learn $(a_{1},b_{1})$, despite the fact that the contribution of $\sigma^{2}$ in Equation (11) in the main text is not purely additive. However, as we see in Figure 20, estimates are uniformly distributed across the permissible space, except for bunching at the upper and lower bounds of $\Lambda$. Note that for numerical and computational stability, we constrain $a_{1}\in(2,50]$ and $b_{1}\in(0.2,50]$ in this example. This contrasts with similarity between replicates visible in Figure 19, and is thus evidence that $(\hat{a}_{1},\hat{b}_{1})$ have no apparent effect on the value of $D(\lambda^{*})$. We should instead set the prior for $\sigma^{2}$ based on external knowledge of the measurement process for $Y$.
Figure 20: Histograms of scaled estimates of $(a_{1}^{*},b_{1}^{*})$ for the settings considered in Section S5.3. Estimates have been scaled to $[0,1]$ for visualisation purposes using the upper and lower limits defined in Table 3.
The regularisation method we employ in the two other examples in the main text is unlikely to assist in estimating $(a_{1},b_{1})$. Promoting a larger mean log marginal standard deviation, with the knowledge $D(\lambda)$ is insensitive to the value of $(a_{1},b_{1})$, would simply pick the largest possible value for $b_{1}^{2}\mathop{/}\left((a_{1}-1)^{2}(a_{1}-2)\right)$, which occurs when $a_{1}$ is at its minimum allowable value and $b_{1}$ its corresponding maximum.
S6 Additional information for the Preece-Baines example
S6.1 Hyperparameter support $\Lambda$
Table 5 contains the upper and lower limits for each hyperparameter, thus defining the feasible region $\Lambda$.
Parameter ($\theta_{q}$) $\mu_{q}$ – Lower $\mu_{q}$ – Upper $\sigma_{q}$ – Lower $\sigma_{q}$ – Upper
$\theta_{1}=h_{0}$ 130 185 $\epsilon$ 30
$\theta_{2}=\delta_{h}$ $\epsilon$ 30 $\epsilon$ 2
$\theta_{3}=s_{0}$ $\epsilon$ 0.2 $\epsilon$ 0.1
$\theta_{4}=\delta_{s}$ $\epsilon$ 1.5 $\epsilon$ 0.2
$\theta_{5}=\gamma$ 9 15 $\epsilon$ 1
Table 5: Parameter vector $\theta$ and associated model specific parameter. The rightmost four columns of the table define the upper and lower limits for the hyperparameters $(\mu_{q},\sigma_{q})$, thus defining $\Lambda$. Informative bounds are required for numerical stability of the data generating process, and an $\epsilon=10^{-6}$ is required to avoid nonsensical optimal values of $\lambda$.
S6.2 Details for $\text{T}(Y)$ and $\text{T}(Y\mid X_{r})$
Denote with $\text{Gamma}(Y;\alpha,\beta)$ the CDF of the gamma distribution with shape parameter $\alpha$ and rate $\beta$; and $\text{Normal}(Y;\xi,\omega^{2})$ the CDF of the normal distribution with mean $\xi$ and standard deviation $\omega$. We define the covariate-independent target
$\text{T}(Y)=0.38\,\text{Gamma}(Y;45.49,0.44)+0.36\,\text{Gamma}(Y;115.41,0.81)+0.27\,\text{Gamma}(Y;277.51,1.64),$ (18)
and the covariate-specific target
$\begin{gathered}\text{T}(Y\mid X_{1}=2)=\text{Normal}(Y;88,3.5^{2}),\quad\text{T}(Y\mid X_{2}=8)=\text{Normal}(Y;130,5.5^{2}),\\ \text{T}(Y\mid X_{3}=13)=\text{Normal}(Y;160,8^{2}),\quad\text{T}(Y\mid X_{4}=18)=\text{Normal}(Y;172,9.5^{2}).\end{gathered}$ (19)
S6.3 Hartmann et al. (2020) priors
Table 6 contains the priors elicited by Hartmann et al. (2020) (extracted from the supplementary material of that paper) for the parameters in the Preece-Baines example. To generate the prior predictive samples displayed in Figure 12 in the main text, we draw, for each user, $\theta$ from the corresponding lognormal distribution then compute $h(t;\theta)$ using (13) (also in the main text, without the error term) at 250 values of $t$ spaced evenly between ages $2$ and $18$.
User Parameter Expectation Variance Lognormal $\mu$ Lognormal $\sigma$
1 $h_{0}$ 162.80 4.20 5.09 0.01
1 $h_{1}$ 174.50 0.80 5.16 0.01
1 $s_{0}$ 0.10 0.10 -3.50 1.55
1 $s_{1}$ 3.30 0.21 1.18 0.14
1 $\theta$ 13.40 0.01 2.60 0.01
2 $h_{0}$ 153.73 1.60 5.04 0.01
2 $h_{1}$ 191.74 4.32 5.26 0.01
2 $s_{0}$ 0.04 0.01 -4.21 1.41
2 $s_{1}$ 2.00 4.30 0.33 0.85
2 $\theta$ 15.90 0.70 2.76 0.05
3 $h_{0}$ 148.80 1.86 5.00 0.01
3 $h_{1}$ 177.14 3.68 5.18 0.01
3 $s_{0}$ 0.07 0.00 -2.75 0.43
3 $s_{1}$ 4.54 37.83 0.99 1.02
3 $\theta$ 11.31 0.21 2.42 0.04
4 $h_{0}$ 162.80 0.02 5.09 0.00
4 $h_{1}$ 174.50 0.01 5.16 0.00
4 $s_{0}$ 0.10 0.01 -2.65 0.83
4 $s_{1}$ 1.60 1.70 0.22 0.71
4 $\theta$ 14.70 0.90 2.69 0.06
5 $h_{0}$ 162.60 0.85 5.09 0.01
5 $h_{1}$ 174.40 0.90 5.16 0.01
5 $s_{0}$ 0.10 0.01 -2.65 0.83
5 $s_{1}$ 3.40 0.01 1.22 0.03
5 $\theta$ 14.60 0.02 2.68 0.01
Table 6: Priors elicited by Hartmann et al. (2020) for each of the 5 users they study. Hartmann et al. provide their results in the form of expected values and variances for the parameters of the model, we compute the corresponding lognormal location $\mu$ and scale $\sigma$ parameters from this information. Values are rounded to two digits of precision.
S6.4 Pareto frontiers for the covariate-independent target
The Pareto frontier for the covariate-independent target and all values of $\kappa\in\mathcal{K}$ is displayed in Figure 21.
[…]
S6.5 Final predictive discrepancy values for the human growth example
The located optimal values of predictive discrepancy are displayed in Figure 22.
Figure 22: Final predictive discrepancy $\log(D(\lambda^{*}))$, or $\log(D(\lambda^{*}\mid\boldsymbol{X}))$ for the covariate-specific target. The multiple objective optimisation approach uses the previously selected values of $\kappa=0.15$ for the covariate-independent target, and $\kappa=0.2$ for the covariate-specific target. Horizontal jitter has been applied for readability.
S6.6 Full marginal prior and posterior comparison plots
Figures 23 and 24 are extended versions of Figure 14 in the main text, and display the prior and posterior estimates for all the parameters in $\theta$. Consistency and uniqueness remain, evidently, challenging and as yet unobtainable.
Figure 23: A comparison of the priors (blue) produced by our method using the covariate-independent marginal target (bottom two rows); and Hartmann et al. (2020) (second row), with no prior displayed for the flat prior scenario. The corresponding posteriors (red) for individual $n=26$ under each of these priors are displayed as dashed lines. Note that y-axes change within columns and are limited to values that clip some of the priors/posteriors for readability.
Figure 24: Otherwise identical to Figure 23 but the bottom two rows display the results obtained using the covariate-specific target.